OpenStudy (anonymous):

write the slope intercept equation for the line that passes through (5,-4) and (7,8)

OpenStudy (jlastino):

first find the slope of the two points (y2-y1)/(x2-x1) x1 = 5 x2= 7 y1= -4 y2=8 The value you get there will be called "m" then use the equation y-y1=m(x-x1) sub the value you got along with the values before

OpenStudy (anonymous):

Thanks so much!!!!!